OpenStudy (anonymous):

Simplify the following expression, and rewrite it in an equivalent form with positive exponents. -30x^4y^-4/120x^-2y^-4

OpenStudy (anonymous):

This is what I got for an answer let me know if I am right (-30/120) * (x^4y^-4/x^2y^-4) -4y/x^2

OpenStudy (anonymous):

jhonyy is that right so far or is it wrong

jhonyy9 (jhonyy9):

is wrong

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

so -30/120 =

OpenStudy (anonymous):

.4

jhonyy9 (jhonyy9):

in farction form

jhonyy9 (jhonyy9):

fraction

jhonyy9 (jhonyy9):

i like yoi learn it

jhonyy9 (jhonyy9):

you

jhonyy9 (jhonyy9):

camon

OpenStudy (anonymous):

in fraction form is 2/5

jhonyy9 (jhonyy9):

why ?

OpenStudy (anonymous):

2 divides into 5 .4 times

jhonyy9 (jhonyy9):

-30/(4*30)

OpenStudy (anonymous):

-225 is the answer for that

jhonyy9 (jhonyy9):

no than will be -1/4

OpenStudy (anonymous):

how?

jhonyy9 (jhonyy9):

so x4/x(-4) = x(4-(-4))

jhonyy9 (jhonyy9):

how ? because 120=4*30

jhonyy9 (jhonyy9):

30 simplify by 30 and will remain -1/4

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

do you understand now sure ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

ok so there is x4/x(-2) than will be x on (4-(-2)) = x on 4+2 = x6

jhonyy9 (jhonyy9):

yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so y(-4)/y(-4) = ?

jhonyy9 (jhonyy9):

a/a = ?

jhonyy9 (jhonyy9):

camon

jhonyy9 (jhonyy9):

easy

OpenStudy (anonymous):

y^16

jhonyy9 (jhonyy9):

no

OpenStudy (anonymous):

y^8

jhonyy9 (jhonyy9):

a/a = ?

OpenStudy (anonymous):

a^2

jhonyy9 (jhonyy9):

a divide a

OpenStudy (anonymous):

=1

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

y^-4/y^-4 = y

jhonyy9 (jhonyy9):

because y(-4)/y(-4) =1 because are the same terms

jhonyy9 (jhonyy9):

is 1

jhonyy9 (jhonyy9):

right

jhonyy9 (jhonyy9):

understandably now ?

OpenStudy (anonymous):

yes but what is my final answer gonna be 1/4x^6/y or turn it around?

jhonyy9 (jhonyy9):

-x6/4

OpenStudy (anonymous):

okay so I am not gonna have the y in there right?

jhonyy9 (jhonyy9):

yes your answer without y

jhonyy9 (jhonyy9):

but with minus

OpenStudy (anonymous):

okay now I understand

jhonyy9 (jhonyy9):

(-1/4)(x6) or (-x6)/4

jhonyy9 (jhonyy9):

understandably sure ?

OpenStudy (anonymous):

yep understanbly

jhonyy9 (jhonyy9):

do you have again ?

jhonyy9 (jhonyy9):

anything

OpenStudy (anonymous):

nope I have a different question now to ask

OpenStudy (anonymous):

Simplify the following expression \[\sqrt{-121} \sqrt{-24}\]

jhonyy9 (jhonyy9):

121=11*11 -1=i*i 24=6*4

jhonyy9 (jhonyy9):

minus * minus = ???

OpenStudy (anonymous):

okay thought that was right but wasn't sure if it had the "i" in it or not

OpenStudy (anonymous):

positive

jhonyy9 (jhonyy9):

yes than result that in your answer not will exist i

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

than sqrt(-121*(-24))= ?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

53.9

jhonyy9 (jhonyy9):

why ? how you have got it ?

jhonyy9 (jhonyy9):

make the calcule here

OpenStudy (anonymous):

well sqrt -121 is 11 and sqrt of -24 is 4.89 rounded up will be 4.9 so times 11*4.9 and get 53.9

jhonyy9 (jhonyy9):

no

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

there is sqrt(-121)*sqrt(-24)=sqrt(-121*(-24))=sqrt(121*24)=sqrt(11*11*4*6) =11*2sqrt6=22sqrt6

OpenStudy (anonymous):

how do you get a sqrt of 6 ?

jhonyy9 (jhonyy9):

24=6*4

jhonyy9 (jhonyy9):

sqrt24=sqrt(6*2*2)=2sqrt6

OpenStudy (anonymous):

sqrt of 6 is 2

jhonyy9 (jhonyy9):

no sqrt of 4 = sqrt 2*2 =2

OpenStudy (anonymous):

okay your right

jhonyy9 (jhonyy9):

now all is understandably ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

have you again ?

OpenStudy (anonymous):

no that is the last problem

jhonyy9 (jhonyy9):

ok bye

jhonyy9 (jhonyy9):

Andrew from Hungary

OpenStudy (anonymous):

Aaron From US

OpenStudy (anonymous):

And Thank you it was also nice meeting you

jhonyy9 (jhonyy9):

nice to meet you

jhonyy9 (jhonyy9):

good luck

OpenStudy (anonymous):

you too

jhonyy9 (jhonyy9):

bye

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

Start with the inner expression, -30x^4, and work outwards. ( ( ( (-30x^4 )y^-4 )/120 )x^-2 )y^-4 =\[-\frac{30 x^4}{\left(\left(y^4 120\right) x^2\right) y^4}= \]\[-\frac{x^2}{4 y^8} \]